# InfoSteer: Steering Information Utility in Language Model Post-Training

### v1

@@ Abstract
Recent advancements in language models (LMs) gradually ushered in an era where post-training is crucial. Yet, post-training approaches such as supervised fine-tuning (SFT) do not guarantee effective use of knowledge acquired during pretraining. We therefore present InfoSteer, a lightweight method that encourages parametric information utilization in LMs during post-training. This is achieved via treating FFN layer as associate key-value memory, and promotes the use of stored memory vectors via forward-pass interventions or regularization during backpropagation. We find this simple guidance during post-training phase delivers consistent performance improvements across diverse model families–including Qwen, Gemma and Llama-spanning over 15 downstream tasks in both ID and OOD evaluations. Beyond performance gains, we also find that steered LMs can adaptively allocate information-placing more emphasis on generating semantically meaningful tokens, while using fewer resources on simple transition ones (e.g., ‘,’ or ‘and’). Our work underscores that vanilla post-training does not fully leverage pre-training potential, and steering LMs in latent representation space offers a promising approach that enhances both performance and interpretability.
@@ 1 Introduction
The contemporary training pipeline for LMs has standardized around a two-stage process: an initial pre-training phase on extensive, web-scale corpora, followed by a post-training phase utilizing smaller, more curated datasets (Ouyang et al., 2022; Touvron et al., 2023; Bai et al., 2023; Mesnard et al., 2024). A considerable body of literature suggests that the fundamental capabilities and knowledge of these models are predominantly instilled during the pre-training stage (Chung et al., 2022; Anil et al., 2023; Muennighoff et al., 2025). Subsequent post-training techniques are often understood as methods to better organize, elicit, or restructure these pre-existing abilities from the base model (Zhou et al., 2023; Rafailov et al., 2024; Guo et al., 2025; Swamy et al., 2025).
Despite this, an open question remains: do post-training methods actually encourage the model to fully utilize the information encoded during pre-training? In many cases, they may not-since models are not explicitly trained or incentivized to retrieve and apply such knowledge [^1] In this work, the terms ‘information’ and ‘knowledge’ are interchangeably used as descriptive language.optimally (Conmy et al., 2023; Chang et al., 2024; Du et al., 2024a). This insufficient use could result in suboptimal performance on downstream tasks, even when relevant knowledge is already stored in the model’s parameters (Bietti et al., 2023; Kim et al., 2025).
We therefore introduce InfoSteer, a prototype designed to encourage more efficient use of a model’s pre-trained knowledge during the post-training phase (Figure 1). Our approach draws on an associative memory view of the Transformer’s feed-forward network (FFN) layers (Geva et al., 2021; Meng et al., 2022a, b; Geva et al., 2023), wherein the first FFN layer acts as a content-dependent key and the second FFN layer as a learned value memory (key-value memory). Consequently, each layer’s output can be viewed as a weighted sum over key-value memory pairs, where each key acts as a coefficient controlling its corresponding memory.
InfoSteer controls the distribution of keys to manage the information-dense memory vectors learned from pretraining. The goal is to promote a high-entropy key distribution, enabling their corresponding memory vectors to be actively engaged during post-training. Specifically, we use two-sided strategies: (i) intervening on memory vectors with low corresponding key values during the forward pass, and (ii) regularizing the entropy of keys in the gradient flow during backpropagation. These methods can be seamlessly incorporated into a vanilla SFT pipeline for information-steered SFT.
Figure 1: Overview of our proposed InfoSteer framework. The interpretation of Transformer FFNs as key-value memory was introduced by Geva et al. (2021), and we provide further details in §3.
@@ 2 Related Work
Model steering represents an emerging paradigm that aims to guide model behavior, focusing on controlled interventions in the model’s internal representation space (Subramani et al., 2022; Zou et al., 2023; Turner et al., 2024; Li et al., 2024; Cyberey and Evans, 2025). Many work has demonstrated the effectiveness of steering methods across various dimensions, including disentangling human-interpretable concepts (Rumelhart et al., 1986), such as linguistic features (e.g., gender, number) (Hewitt and Manning, 2019; Lasri et al., 2022; Wang et al., 2022; Hanna et al., 2023; Huang et al., 2024; Chang et al., 2025) and logical reasoning (Wu et al., 2024b; Xie et al., 2025). These approaches typically operate by introducing auxiliary objectives, modifying activation patterns, or implementing controlled perturbations during forward or backward passes (Rimsky et al., 2023; Scalena et al., 2024; Stolfo et al., 2024; Luo et al., 2025; Gur-Arieh et al., 2025; Bartoszcze et al., 2025). The key advantage of model steering lies in its ability to leverage the rich knowledge already encoded in pretrained weights while directing how this information is accessed and applied to downstream tasks (Geiger et al., 2023; Lee et al., 2024; Siddique et al., 2025; Soo et al., 2025; Deng et al., 2025). Our information-steered approach extends this paradigm by targeting the key distribution in transformer FFN layers, which influences which memory vectors are activated during computation.
@@
From an information-theoretic perspective, quantifying the information encoded in a LMs’ high-dimensional parameters space remains a challenging problem (Achille et al., 2020; Bernstein and Yue, 2021). This challenge is closely tied to the ongoing discourse on parametric versus non-parametric knowledge storage in LMs (Tay et al., 2022; Ferrando et al., 2022, 2023; Xie et al., 2023; Deng et al., 2024; Ferrando and Voita, 2024; Du et al., 2024b) (or token-mixing in attn layer vs. channel mixing in FFN layer). The concept of channel mixing stems from the perspective of viewing the FFN layer as a key-value memory structure (Weston et al., 2014; Geva et al., 2021; Qiu et al., 2024; Kim et al., 2025) . Empirical evidence supports this interpretation, showing that individual neurons in FFN layers activate in response to specific semantic concepts and linguistic patterns (Dai et al., 2021; Liu et al., 2023; Niu et al., 2024). Furthermore, intervention studies have demonstrated that targeted modifications to these memory vectors can directly influence model outputs on knowledge-intensive tasks (Meng et al., 2022a, b; Geva et al., 2023; Hase et al., 2023; Yao et al., 2024; Wang et al., 2024), suggesting that factual information is strongly correlated with in this component.
@@ 3 Transformer Feed-Forward Layers as Key-Value Memories
$\text{FFN}(h)=\sum_{i=1}^{d_{m}}\underbrace{\sigma(hW_{\text{up}})_{i}}_{\text{key coefficient}}\cdot\underbrace{(W_{\text{down}})_{i,:}}_{\text{value vector}}=k_{1}v_{1}+k_{2}v_{2}+\cdots+k_{d_{m}}v_{d_{m}}$ (3)
where each $k_{i}\in\mathbb{R}$ is a scalar key coefficient, and each $v_{i}\in\mathbb{R}^{d}$ is a value vector (memory vector). The key coefficients control the magnitude of the memory vectors’ contribution to the final prediction.
@@ 4 InfoSteer
LMs are commonly interpreted to store parametric knowledge within their dense layers, particularly in the FFN blocks. (Meng et al., 2022a; Geva et al., 2023; Kim et al., 2025). However, during post-training, there is no explicit guidance to encourage the model to utilize this knowledge for new alignment or retrieval tasks. InfoSteer is proposed to bridge this gap.
4.1 Key Motivation & Desierata
To steer the information utility of pretrained LMs, we aim to enhance the engagement of memory vectors during post-training. The core idea is to control the distribution of key coefficients in the position-wise FFN (see Figure 2). Specifically, we observe a dominant property of associative memory: if a key coefficient $k_{i}$ is significantly larger than another key coefficient $k_{j}$, the final prediction predominantly relies on the corresponding memory vector $v_{i}$, while $v_{j}$ is underutilized or overlooked. Our goal is to achieve two objectives:
@@ 4.1 Key Motivation & Desierata
    Minimize the language modeling loss: Ensure the FFN output $\text{FFN}(h)=\sum_{i=1}^{d_{m}}k_{i}v_{i}$ contributes to accurate predictions.
@@
Promoting entropy weaken the dominance of individual coefficients, encouraging more balanced activation of memory vectors $v_{i}$ and thus enabling richer parametric knowledge retrieval.
4.2 Methods: Intervention and Regularization
@@
$k_{i}^{(l)}\leftarrow\alpha\cdot\frac{1}{d_{m}}\sum_{i=1}^{d_{m}}k_{i}^{(l)},\text{for}\enspace i\in\mathcal{I}^{(l)}.$ (4)
This adjustment ensures that previously minor coefficients contribute more significantly to the FFN output $\text{FFN}(h)=\sum_{i=1}^{d_{m}}k_{i}v_{i}$. $\alpha$ is controlling scalar to set the magnitude of steering.
@@ 4.3 MLP Variants
When adapting methods for MLP variants like SwiGLU, used in models such as Qwen (Bai et al., 2023), LLaMA (Touvron et al., 2023), and Gemma (Mesnard et al., 2024), we focus on the input to the down projection. SwiGLU modifies the standard FFN as:
@@
where $\sigma$ is the silu activation, and $\odot$ denotes the elment-wise multiplication. Similar to standard two-layer FFN, we modulate the key coefficient distribution $\mathbf{k}=\sigma(hW_{\text{gate}})\odot(hW_{\text{up}})$.
@@ 5 Experiment
To evaluate the overall effectiveness of steered SFT against vanilla SFT, we conduct a comprehensive study across various downstream tasks and model sizes. Our primary focus is on achieving a holistic understanding of the overall model behavior after applying our methods, with particular attention paid to generalization performance. The details are provided below [^1] In this paper, we primarily focus on a generic method. But we also provide fine-grained steering methods in Appendix B and the ablation study of different layers in Appendix C..
@@ 5.1 Experiment Setup
We use standard SFT as our baseline. For information-steered SFT, we evaluate intervention and regularization methods, which share the goal of enhancing memory engagement during training. The default hyperparameters for the intervention are set to $p$= 0.01 and $\alpha=1$, with $\lambda=0.01$ for entropy regularization. Other training details are provided in Appendix A.
@@ 5.2 General Performance Comparison
The benefits of steering reveal a critical limitation in contemporary large language models: they substantially underutilize the knowledge acquired during pretraining when applied to downstream tasks. This inefficiency is strikingly evident across all three widely-used model families examined-Gemma, LLaMA, and Qwen. Specifically, the Gemma-2-9B base model achieves $90.1$% on HellaSwag, yet reaches $95.7$% with intervention steering-a $5.6$% improvement without additional pretraining or parameter increase. Similarly, LLaMA-3-8B shows a remarkable $5.5$% gain on HellaSwag with intervention steering over the base model. This pattern repeats across all architectures and benchmarks, suggesting that modern LLMs possess substantially more capabilities than they successfully deploy in standard fine-tuning regimes.

### v2

# Steering Information Utility in Key-Value Memory for Language Model Post-Training

@@ Abstract
Recent advancements in language models (LMs) have marked a shift toward the growing importance of post-training. Yet, post-training approaches such as supervised fine-tuning (SFT) do not guarantee the effective use of knowledge acquired during pretraining. We therefore introduce InfoSteer, a lightweight method that encourages parametric information utilization in LMs during post-training. Specifically, InfoSteer treats the feed-forward network (FFN) layer as associate key-value memory and promotes the use of stored memory vectors via forward-pass interventions or regularization during backpropagation. This simple guidance during post-training phase yields consistent performance improvements across diverse model families–including Qwen, Gemma and Llama—spanning 15 downstream tasks in both in-distribution (ID) and out-of-distribution (OOD) evaluations. Beyond performance gains, we also find that steered LMs can adaptively allocate information by placing more emphasis on generating semantically meaningful tokens, while using fewer resources on simple transition ones (e.g., ‘,’ or ‘and’). Our work underscores that vanilla post-training does not fully exploit the potential gained during pre-training, and that steering LMs in latent representation space offers a promising approach to enhance both performance and interpretability. [^1] The code is available at: https://github.com/chili-lab/InfoSteer.
@@ 1 Introduction
The contemporary training pipeline for LMs has standardized around a two-stage process: an initial pre-training phase on extensive, web-scale corpora, followed by a post-training phase utilizing smaller, more curated datasets (Ouyang et al., 2022; Touvron et al., 2023; Bai et al., 2023; Mesnard et al., 2024). A considerable body of literature suggests that the fundamental capabilities and knowledge of these models are predominantly instilled during the pre-training stage (Chung et al., 2022; Anil et al., 2023; Muennighoff et al., 2025). Subsequent post-training techniques are commonly viewed as approaches to better refine, elicit, or adapt these inherent capabilities embedded in the base model (Zhou et al., 2023; Rafailov et al., 2024; Guo et al., 2025; Swamy et al., 2025).
Despite this, an open question remains: do post-training methods truly encourage the model to fully utilize the information encoded during pre-training? In many cases, they may not–since models are neither explicitly trained nor incentivized to retrieve and apply such knowledge [^2] In this work, the terms ‘information’ and ‘knowledge’ are interchangeably used as descriptive language. optimally (Conmy et al., 2023; Chang et al., 2024; Du et al., 2024a). This insufficient use could result in suboptimal performance on downstream tasks, even when relevant knowledge is already stored in the model’s parameters (Bietti et al., 2023; Kim et al., 2025).
We therefore introduce InfoSteer, a method designed to encourage more effective use of a model’s pre-trained knowledge during the post-training phase (Figure 1). Our approach draws on an associative memory view of the Transformer’s feed-forward network (FFN) layers (Geva et al., 2021; Meng et al., 2022a, b; Geva et al., 2023), where the first FFN layer acts as a content-dependent key and the second FFN layer functions as a learned value memory, together forming a key–value memory mechanism. Consequently, each layer’s output can be viewed as a weighted sum of memory vectors, where each vector’s weight (the key coefficient) is produced by the input vector multiplied by its corresponding key in the first FFN. [^3] In this work, we term the output of first FFN as key coefficient and its distribution as key distribution. This is the same as “memory coefficient” in  Geva et al. (2021)
InfoSteer controls the distribution of key coefficients to manage the information-dense memory vectors learned from pretraining. The goal is to promote a high-entropy key distribution, enabling their corresponding memory vectors to be actively engaged during post-training. Specifically, we use two strategies: (i) intervening on memory vectors with low corresponding key coefficients during the forward pass, and (ii) regularizing the entropy of key coefficients in the gradient flow during backpropagation. These methods can be seamlessly incorporated into a vanilla SFT pipeline for information-steered SFT.
Figure 1: Overview of our proposed InfoSteer framework. The interpretation of Transformer FFNs as key-value memory was introduced by Geva et al. (2021), further details are provided in §3.
@@ 2 Related Work
Model steering represents an emerging paradigm to guide model behavior, focusing on controlled interventions in the model’s latent space (Subramani et al., 2022; Zou et al., 2023; Turner et al., 2024; Li et al., 2024; Cyberey and Evans, 2025). Many work has demonstrated the effectiveness of steering methods across various dimensions, including disentangling human-interpretable concepts (Rumelhart et al., 1986), such as linguistic features (e.g., gender, number) (Hewitt and Manning, 2019; Lasri et al., 2022; Wang et al., 2022; Hanna et al., 2023; Huang et al., 2024; Chang et al., 2025) and logical reasoning (Wu et al., 2024b; Xie et al., 2025). These approaches typically operate by introducing auxiliary objectives, modifying activation patterns, or implementing controlled perturbations during forward or backward passes (Rimsky et al., 2023; Scalena et al., 2024; Stolfo et al., 2024; Luo et al., 2025; Gur-Arieh et al., 2025; Bartoszcze et al., 2025). The key advantage of model steering lies in its ability to leverage the rich knowledge already encoded in pretrained weights while directing how this information is accessed and applied to downstream tasks (Geiger et al., 2023; Lee et al., 2024; Siddique et al., 2025; Soo et al., 2025; Deng et al., 2025). Our information-steered approach extends this paradigm by targeting the key distribution in transformer FFN layers, which influences which memory vectors are activated during computation.
@@
From an information-theoretic perspective, quantifying the information encoded in an LM’s high-dimensional parameter space remains a challenging problem (Achille et al., 2020; Bernstein and Yue, 2021). This challenge is closely tied to the ongoing discourse on parametric versus non-parametric knowledge storage in LMs (Tay et al., 2022; Ferrando et al., 2022, 2023; Xie et al., 2023; Deng et al., 2024; Ferrando and Voita, 2024; Du et al., 2024b) (or token-mixing in attention layer vs. channel mixing in FFN layer). The concept of channel mixing stems from the perspective of viewing the FFN layer as a key-value memory structure (Weston et al., 2014; Geva et al., 2021; Qiu et al., 2024; Kim et al., 2025) . Empirical evidence supports this interpretation, showing that individual neurons in FFN layers activate in response to specific semantic concepts and linguistic patterns (Dai et al., 2021; Liu et al., 2023; Niu et al., 2024). Furthermore, intervention studies have demonstrated that targeted modifications to these memory vectors can directly influence model outputs on knowledge-intensive tasks (Meng et al., 2022a, b; Geva et al., 2023; Hase et al., 2023; Yao et al., 2024; Wang et al., 2024), suggesting that factual information is strongly correlated with in this component.
@@ 3 Transformer Feed-Forward Layers as Key-Value Memories
$\text{FFN}(h)=\sum_{i=1}^{d_{m}}\underbrace{\sigma(hW_{\text{up}})_{i}}_{\text{key coefficient}}\cdot\underbrace{(W_{\text{down}})_{i,:}}_{\text{value vector}}=k_{1}\mathbf{v}_{1}+k_{2}\mathbf{v}_{2}+\cdots+k_{d_{m}}\mathbf{v}_{d_{m}}$ (3)
where each $k_{i}\in\mathbb{R}$ is a scalar key coefficient, and each $\mathbf{v}_{i}\in\mathbb{R}^{d}$ is a value vector (memory vector). The key coefficients control the magnitude of the memory vectors’ contribution to the final prediction.
@@ 4 InfoSteer
LMs are commonly interpreted to store parametric knowledge within their dense layers, particularly in the FFN blocks (Meng et al., 2022a; Geva et al., 2023; Kim et al., 2025). However, during post-training, there is no explicit guidance to encourage the model to utilize this knowledge for new alignment or retrieval tasks. Therefore, we propose InfoSteer to bridge this gap.
4.1 Motivation & Desierata
To steer the information utility of pretrained LMs, we aim to enhance the engagement of memory vectors during post-training. The core idea is to control the distribution of key coefficients in the position-wise FFN (see Figure 2). Specifically, we observe a dominant property of associative memory: if a key coefficient $k_{i}$ is significantly larger than another key coefficient $k_{j}$, the final prediction predominantly relies on the corresponding memory vector $\mathbf{v}_{i}$, while $\mathbf{v}_{j}$ is underutilized or overlooked. Our goal is to achieve two objectives:
@@ 4.1 Motivation & Desierata
    Minimize the language modeling loss: Ensure the FFN output $\text{FFN}(h)=\sum_{i=1}^{d_{m}}k_{i}\mathbf{v}_{i}$ contributes to accurate predictions.
@@
Promoting entropy weaken the dominance of individual coefficients, encouraging more balanced activation of memory vectors $\mathbf{v}_{i}$ and thus enabling richer parametric knowledge retrieval.
4.2 Generic Methods: Intervention and Regularization
@@
$k_{s}^{(l)}\leftarrow\alpha\cdot\frac{1}{d_{m}}\sum_{i=1}^{d_{m}}k_{i}^{(l)},\text{for}\kern 5.0pts\in\mathcal{I}^{(l)}.$ (4)
This adjustment ensures that previously minor coefficients contribute more significantly to the FFN output $\text{FFN}(h)=\sum_{i=1}^{d_{m}}k_{i}\mathbf{v}_{i}$. $\alpha$ is a controlling scalar to set the magnitude of steering.
@@
4.3 Fine-Grained Steering of Memory Vectors
While the generic methods modulate the overall key distribution, a more targeted approach requires first understanding what information each memory vector $\mathbf{v}_{i}$ encodes. The information surrogate $\phi_{i}$ provides a direct, formalized method for this characterization.
Let $v_{i}\in\mathbb{R}^{d}$ be the $i$-th memory vector (i.e., the $i$-th row of $W_{\text{down}}$) and $W_{\text{decode}}\in\mathbb{R}^{d\times|V|}$ be the language model’s decoding head, where $|V|$ is the vocabulary size. The information surrogate $\phi_{i}\in\mathbb{R}^{|V|}$ for vector $v_{i}$ is the resulting logit vector:
$\phi_{i}=v_{i}\cdot W_{\text{decode}}.$ (6)
This surrogate $\phi_{i}$ acts as a “semantic fingerprint” for $v_{i}$. We propose the following algorithm to characterize $v_{i}$ by analyzing the properties of this surrogate logit vector.
1. Surrogate Normalization.
First, we compute the full probability distribution $P_{i}\in\mathbb{R}^{|V|}$ over the vocabulary by applying the Softmax function to the logit vector $\phi_{i}$, $P_{i}=\text{Softmax}(\phi_{i})$, where the $j$-th element of $P_{i}$, $p_{i,j}=\frac{\exp(\phi_{i,j})}{\sum_{k=1}^{|V|}\exp(\phi_{i,k})}$.
2. Specificity Analysis.
We quantify the vector’s “specificity” by calculating the entropy $H(P_{i})$ of its distribution. This measures the concentration of the distribution $H(P_{i})=-\sum_{j=1}^{|V|}p_{i,j}\log p_{i,j}$. A low $H(P_{i})$ indicates a high-specificity vector, as its probability mass is concentrated on a narrow set of tokens (representing specialized knowledge) and vice versa.
3. Semantic Concept Identification.
For vectors identified as specialized (e.g., $H(P_{i})<\tau$ for some entropy threshold $\tau$), we identify their semantic focus. We find the set of indices $\mathcal{I}_{i}$ corresponding to the $K$ largest probabilities in $P_{i}$:
$\mathcal{I}_{i}=\underset{j\in\{1,\dots,|V|\}}{\text{arg top-}K}\,(p_{i,j}).$ (7)
The semantic concept, $T_{i}$, is the set of vocabulary tokens corresponding to these indices:
$T_{i}=\{\text{token}(j)\mid j\in\mathcal{I}_{i}\}.$ (8)
This set $T_{i}$ (e.g., {‘quantum’, ‘physics’, ‘superposition’}) reveals the specific topic or concept encoded by the memory vector $v_{i}$.
The results of this characterization—such as the entropy $H(P_{i})$ and the semantic concept set $T_{i}$ for each vector $v_{i}$—can then be used to inform a more targeted intervention or regularization strategy. For example, one could selectively amplify key coefficients $k_{i}$ corresponding to vectors $v_{i}$ whose $T_{i}$ matches a desired topic. [^4] In this paper, we primarily focus on a generic method, which provides substantial performance gains. While fine-grained steering is a promising approach for encouraging LMs to utilize specific knowledge during post-training, our investigation found it offered only marginal improvements over the generic method. Therefore, we detail these fine-grained methods in Appendix B and provide a layer-wise ablation study in Appendix C, leaving more granular control of knowledge steering as a direction for future research.
@@ 4.4 MLP Variants
Modern LM architectures may include MLP variants that differ from the typical FFN. For example, when adapting methods for MLP variants like SwiGLU, used in models such as Qwen (Bai et al., 2023), LLaMA (Touvron et al., 2023), and Gemma (Mesnard et al., 2024), we focus on the input to the down projection. SwiGLU modifies the standard FFN as:
@@
where $\sigma$ is the SiLU activation, and $\odot$ denotes element-wise multiplication.
Similar to a standard two-layer FFN, we define the key coefficient distribution as $\mathbf{k}=\sigma(hW_{\text{gate}})\odot(hW_{\text{up}})$. The core idea for identifying this key distribution is that it represents the input just before it associates with the memory vectors (i.e., $W_{\text{down}}$). Regardless of whether a gated function or another method is used to derive it, our focus remains on this input $\mathbf{k}$, as it is what directly manipulates the memory vectors.
@@ 5 Experiment
To evaluate the overall effectiveness of steered SFT against vanilla SFT, we conduct a comprehensive study across various downstream tasks and model sizes. Our primary focus is on achieving a holistic understanding of the overall model behavior after applying our methods, with particular attention paid to generalization performance. The details are provided below.
@@ 5.1 Experiment Setup
We use standard SFT as our baseline. For information-steered SFT, we evaluate intervention and regularization methods, which share the goal of enhancing memory engagement during training. The default hyperparameters for the intervention are set to pp% = 0.01 and $\alpha=1$, with $\lambda=0.01$ for entropy regularization. Other training details are provided in Appendix A.
@@ 5.2 General Performance Comparison
The benefits of steering reveal a critical limitation in contemporary large language models: they substantially underutilize the knowledge acquired during pretraining when applied to downstream tasks. This inefficiency is strikingly evident across all three widely-used model families examined-Gemma, LLaMA, and Qwen. Specifically, the Gemma-2-9B base model achieves 90.190.1% on HellaSwag, yet reaches 95.795.7% with intervention steering-a 5.65.6% improvement without additional pretraining or parameter increase. Similarly, LLaMA-3-8B shows a remarkable 5.55.5% gain on HellaSwag with intervention steering over the base model. This pattern is consistent across all architectures and benchmarks, suggesting that modern LLMs possess substantially more capabilities than they successfully deploy in standard fine-tuning regimes.
Table 1: Performance comparison of Qwen, LLaMA, and Gemma models with different training methods on eight datasets. To highlight improvements, we use blue to highlight significant gains. All results are reported as the average scores over three independent runs.